Complete the following: Magnitude Number 4.30 A 2.5x -3.5y Number Number Number r + Number Number Number r + Number Number Number r +
Answer #1

As2.5i-3.5ỳ magnitude-A-V2.52+3.5.14.301] 2xA-5.0x-7ný magnitude =V52 + 72 =8.6021 -2 XÃ--50x + 7.0 magnitude |(-5)-+ 7-8602 magnitude A2.5 3.54.301 =1.250 x + 1.750 2 / 1.2502+1.75% |2.151 ] magnitude